OpenStudy (anonymous):

how do i find xf knowing xi= 0 m/s, a= -5.31 m/s^2, delta t= 7.5s

OpenStudy (anonymous):

\[v=u+a \Delta t\]

OpenStudy (anonymous):

is this how to find the displacement?

OpenStudy (anonymous):

" xi= 0 m/s " did you means Vi = 0 m/s or xi = 0m either way the equation is xf = xi +Vi*t - (a*t^2)/2